From a point on a straight​ road, PabloPablo and MariaMaria ride bicycles in the same direction. PabloPablo rides at 77 mph and MariaMaria rides at 1111 mph. In how many hours will they be 3434 mi​ apart?
Found 2 solutions by jorel555, MathTherapy:
Answer by jorel555(1290)

If, at the end of time t, MariaMaria and PabloPablo are 3434 miles apart, then:
77t+3434=1111t
1034t=3434
t=3434/1034=3.32 hours of riding until they are the required distance apart!!!!!!!

Answer by MathTherapy(10555)


From a point on a straight​ road, PabloPablo and MariaMaria ride bicycles in the same direction. PabloPablo rides at 77 mph and MariaMaria rides at 1111 mph. In how many hours will they be 3434 mi​ apart?
The numbers make ABSOLUTELY NO SENSE. If you really need help, don't you think you should at least post correct info.?
What is wrong with some of you? Don't you realize how unrealistic and senseless this post is?